# Towards Visual Text Grounding of Multimodal Large Language Model

### v1

@@ 5.2 Further Findings
Our settings reflect both MLLMs’ instruction-following and grounding ability. Most of the existing open-source models are not able to follow the relatively complex instructions [30] in this task, and even if they correctly follow the instructions, most of them are not able to find the correct grounded bounding boxes that support their answers. Comparing these results with their outstanding performances on standard QA tasks, we reveal this critical issue that most existing open-source MLLMs are potentially overfitted to the standardized tasks, and they still lack instruction-following abilities.
@@ 5.3 Ablation Study
Effect of Embedding Merging. We evaluate the effectiveness of utilizing our proposed embedding merging techniques. As shown in Table 4, without embedding merging, the average performance is the lowest among the settings in which the embedding merging is utilized. In the table, $3\times 3$ Similarity Merging represents merging the similarities within the $3\times 3$ surrounding window of each patch. Among all the experimental settings, the utilization of $3\times 3$ embedding merging reaches the best performance. The results are reasonable since embedding merging with a smaller window size can alleviate the randomness while a larger window might smooth the values too much, causing the similarities to be too closed to distinguish.
@@ 6.1 Practical Use Case
This text-rich document grounding capability helps users understand the source of the information provided in the answers, building users’ trust in the system. When users ask questions to AI assistants about documents, they need to verify the answer. Without grounding, they may look through the whole document, which damages the user experience. However, if grounded boxes are given, users can verify the answer with only a glimpse, largely accelerating users’ reliance on AI assistants for their daily work. Besides, text grounding has been deployed in real-world products, while visual text grounding remains challenging. Thus, this benchmark is practically useful.
@@ 6.2 Differences Between Common Text Localization Task
Our task focuses on text grounding in document images, which is different from text localization and orthogonal to all existing tasks: Text localization requires the model to provide the location of specific visual text or extract from the page. For example, the text localization used in KOSMOS-2.5 [46] contains Document-Level Text Recognition and Image-to-Markdown Generation, which are totally different from ours. On the contrary, our visual text grounding aims to provide the answer with its sources from the image. The key difference lies in that our prediction of text position is coupled with the understanding of image content.
@@ 7 Conclusion
In this paper, we introduce TRIG, a novel task designed to evaluate and enhance the visual text grounding capabilities of MLLMs in text-rich document images. We first build TRIG-Bench, a comprehensive benchmark and a large-scale instruction tuning dataset. Our evaluation reveals significant limitations in the visual text grounding capabilities of existing MLLMs, particularly in their ability to handle complex document layouts and textual content. Hence, we propose two simple and effective instruction-tuning-based and embedding-based methods that demonstrate substantial improvements over existing models. We believe this task is still an under-explored task that has not been well solved and might require more attention from the community.
@@ Appendix B Training Configurations
Our training data contains $22,363$ samples from CharQA, $29,094$ samples from DocVQA, $19,840$ samples from InfographisVQA, and $24,885$ samples from TRINS. All experiments are performed on Nvidia A100 GPUs. We utilize $8$ GPUs for training and $1$ for inference.
@@
For the embedding-based method, we utilize PaliGemma-3B [3] as our base model to be finetuned, based on the codebase from ColPali [14]. We train our model for $4$ epochs with a batch size of $32$ and a learning rate of $55e-$ 5, using the AdamW optimizer. LoRA adapters with a rank of $32$ are added to all MLP layers in the LLM.
@@ C.2 Evaluation Setting 2 (OCR-based Grounding)
[…]
@@ E.2 Evaluation Prompts
In order to alleviate the influences of prompts when testing on our benchmark, 3 different prompts with diverse formatting requirements [39, 30, 27] are proposed for each evaluation setting, resulting in a total of 9 different evaluation prompts. For each evaluation setting, we present the best results across the 3 different prompts as shown in Figure 9. The main difference between different prompts in the same evaluation setting is the required format for the generation or selection of the grounded bounding boxes. Specifically, prompts with “#1” utilize the CSS format, which is widely used for pretraining; prompts with “#2” utilize the most naive format containing the necessary information in a list; prompts with “#3” are similar to “#2” but utilize the relative coordinates.

### v2

@@ 5.2 Further Findings
Our settings reflect both MLLMs’ instruction-following and grounding ability. Most of the existing open-source models are not able to follow the relatively complex instructions in this task, and even if they correctly follow the instructions, most of them are not able to find the correct grounded bounding boxes that support their answers. Comparing these results with their outstanding performances on standard QA tasks, we reveal this critical issue that most existing open-source MLLMs are potentially overfitted to the standardized tasks, and they still lack instruction-following abilities.
@@ 5.3 Ablation Study
Effect of Embedding Merging. We evaluate the effectiveness of utilizing our proposed embedding merging techniques. As shown in Table 4, without embedding merging, the average performance is the lowest among the settings in which the embedding merging is utilized. In the table, $3\times 3$ Similarity Merging represents merging the similarities within the $3\times 3$ surrounding window of each patch. Among all the experimental settings, the utilization of $3\times 3$ embedding merging reaches the best performance. The results are reasonable since embedding merging with a smaller window size can alleviate the randomness, while a larger window might smooth the values too much, causing the similarities to be too close to distinguish.
@@ 6.1 Practical Use Case
This text-rich document’s grounding capability helps users understand the source of the information provided in the answers, building users’ trust in the system. When users ask questions to AI assistants about documents, they need to verify the answer. Without grounding, they may look through the whole document, which damages the user experience. However, if grounded boxes are given, users can verify the answer with only a glimpse, largely accelerating users’ reliance on AI assistants for their daily work. Besides, text grounding has been deployed in real-world products, while visual text grounding remains challenging. Thus, this benchmark is practically useful.
@@ 6.2 Differences Between Common Text Localization Task
Our task focuses on text grounding in document images, which is different from text localization and orthogonal to all existing tasks: Text localization requires the model to provide the location of specific visual text or extract it from the page. For example, the text localization used in KOSMOS-2.5 [46] contains Document-Level Text Recognition and Image-to-Markdown Generation, which are totally different from ours. On the contrary, our visual text grounding aims to provide the answer with its sources from the image. The key difference lies in that our prediction of text position is coupled with the understanding of image content.
@@ 7 Conclusion
In this paper, we introduce TRIG, a novel task designed to evaluate and enhance the visual text grounding capabilities of MLLMs in text-rich document images. We first built TRIG-Bench, a comprehensive benchmark and a large-scale instruction tuning dataset. Our evaluation reveals significant limitations in the visual text grounding capabilities of existing MLLMs, particularly in their ability to handle complex document layouts and textual content. Hence, we propose two simple and effective instruction-tuning-based and embedding-based methods that demonstrate substantial improvements over existing models.
@@ Appendix B Training Configurations
Our training data contains $22,363$ samples from CharQA, $29,094$ samples from DocVQA, $19,840$ samples from InfographisVQA, and $24,885$ samples from TRINS. All experiments are performed on Nvidia A100 GPUs. We utilize $8$ GPUs for training and 11 for inference.
@@
For the embedding-based method, we utilize PaliGemma-3B [4] as our base model to be finetuned, based on the codebase from ColPali [17]. We train our model for $4$ epochs with a batch size of $32$ and a learning rate of 5e55e-5, using the AdamW optimizer. LoRA adapters with a rank of $32$ are added to all MLP layers in the LLM.
@@ C.2 Evaluation Setting 2 (OCR-based Grounding)
$\text{F1}=\frac{1}{N}\sum_{i=1}^{N}\frac{2\times\text{Precision}_{i}\times\text{Recall}_{i}}{\text{Precision}_{i}+\text{Recall}_{i}}$ (11)
@@ C.3 Extra Evaluation Setting 3
Qwen2-2B 5.35 7.37 11.34 7.69 2.18 2.23 8.75 3.48 0.57 0.75 1.21 0.88 4.09 4.90 6.75 5.05 4.21
Qwen2-7B 3.78 6.08 5.50 4.98 2.17 2.50 2.17 2.25 0.54 0.92 1.00 0.83 2.74 3.91 4.79 3.58 2.95
Qwen2.5-3B 2.46 3.30 6.88 3.86 1.16 1.70 2.50 1.68 0.28 0.38 0.67 0.48 10.44 13.44 19.55 14.07 5.17
Qwen2.5-7B 3.23 5.24 4.91 4.48 5.60 6.08 7.92 6.45 0.37 0.58 0.58 0.58 30.30 36.80 45.18 37.31 12.89
llava-ov-1B 3.84 4.58 12.55 6.31 3.80 3.93 13.46 5.89 1.18 1.50 2.92 1.93 19.34 20.39 38.08 24.59 9.53
llava-ov-7B 0.13 0.25 0.17 0.20 0.08 0.10 0.25 0.14 0.00 0.00 0.00 0.00 0.60 0.63 0.75 0.67 0.25
InternVL2-1B 3.81 3.75 22.88 5.53 0.47 0.45 6.00 0.84 2.27 2.13 16.73 3.66 21.08 22.87 55.22 28.11 9.54
InternVL2-8B 0.23 0.23 2.25 0.42 2.07 2.25 7.58 2.68 0.20 0.22 0.71 0.32 2.59 2.67 5.00 3.23 1.97
InternVL2.5-1B 2.73 2.73 18.58 4.63 0.34 0.34 4.67 0.61 0.15 0.17 1.38 0.29 10.39 14.43 24.47 14.25 5.49
InternVL2.5-8B 3.31 3.88 20.46 5.50 2.29 2.54 10.25 10.25 0.80 0.99 5.21 5.21 24.43 35.36 37.67 30.39 12.96
@@ E.2 Evaluation Prompts
In order to alleviate the influences of prompts when testing on our benchmark, 3 different prompts with diverse formatting requirements [40, 32, 29, 62] are proposed for each evaluation setting, resulting in a total of 9 different evaluation prompts. For each evaluation setting, we present the best results across the 3 different prompts as shown in Figure 9. The main difference between different prompts in the same evaluation setting is the required format for the generation or selection of the grounded bounding boxes. Specifically, prompts with “#1” utilize the CSS format, which is widely used for pretraining; prompts with “#2” utilize the most naive format containing the necessary information in a list; prompts with “#3” are similar to “#2” but utilize the relative coordinates.